# The large-$N$ limit of the topological susceptibility of $\mathrm{SU}(N)$ Yang–Mills theories via Parallel Tempering on Boundary Conditions

### v1

@@ 4 Conclusions
I have presented results for the autocorrelation times of PTBC and discussed their scaling with the lattice spacing and with the number of colors, showing the improvement achieved with respect to standard algorithms. The adoption of PTBC permitted me to determine the topological susceptibility with high accuracy for finer lattice spacings compared to previous investigation with periodic and open boundary conditions, allowing me to achieve controlled continuum limits of $\chi$ for all $N$. By taking the continuum limit of $\chi$ at fixed smoothing radius in physical units for several choices of $R_{\scriptscriptstyle{\mathrm{s}}}$, I also showed the independence of $\chi$ on $R_{\scriptscriptstyle{\mathrm{s}}}$ in the continuum limit in the probed range of smoothing radii. Combining my extrapolated result for the large-$N$ limit of $\chi$ with a recent large-$N$ determination of $F_{\pi}$, I have found excellent agreement both with the Witten–Veneziano formula, and with the Di Vecchia–Veneziano equation describing the relation between quenched and unquenched $\chi$ at large $N$. Finally, I have performed a comprehensive comparison of my determinations with previous ones in the literature, generally finding very good agreement with previous PBC and OBC calculations, and clarifying the origin of a few tensions with Ref. Athenodorou:2021qvs ’s results. The most important results presented in this manuscript are summarized in the text box below:

### v2

@@ 4 Conclusions
I have presented results for the autocorrelation times of PTBC and discussed their scaling with the lattice spacing and with the number of colors, showing the improvement achieved with respect to standard algorithms. The adoption of PTBC permitted me to determine the topological susceptibility with high accuracy for finer lattice spacings compared to previous investigations with periodic and open boundary conditions, allowing me to achieve controlled continuum limits of $\chi$ for all $N$. By taking the continuum limit of $\chi$ at fixed smoothing radius in physical units for several choices of $R_{\scriptscriptstyle{\mathrm{s}}}$, I also showed the independence of $\chi$ on $R_{\scriptscriptstyle{\mathrm{s}}}$ in the continuum limit in the probed range of smoothing radii. Combining my extrapolated result for the large-$N$ limit of $\chi$ with a recent large-$N$ determination of $F_{\pi}$, I have found excellent agreement both with the Witten–Veneziano formula, and with the Di Vecchia–Veneziano equation describing the relation between quenched and unquenched $\chi$ at large $N$. Finally, I have performed a comprehensive comparison of my determinations with previous ones in the literature, generally finding very good agreement with previous PBC and OBC calculations, and clarifying the origin of a few tensions with Ref. Athenodorou:2021qvs ’s results. The most important results presented in this manuscript are summarized in the text box below: